I'm going tp put the 4 problems I'm having a hard time with together:
3x^2+20x-3=0
x=?

e=1/7 j(c-d)
c=?

i^18

find the slope of the line containing the pair of points (-6,-9)(12,2)

Answer by stanbon(75887)

3x^2+20x-3=0
Use the quadratic formula:
x=[-20+-sqrt(400+36)]/6=[-20+-20.88]/6=0.467 or -6.81333...

e=1/7 j(c-d)
Do you mean e=(1/7)j(c-d)???
c=?
i^18
=[i^2]^9
=(-1)^9=-1
find the slope of the line containing the pair of points (-6,-9)(12,2)
slope=[-9-2]/[-6-12]= -11/-18=11/18
Cheers,
Stan H.